# Mathematical Language Models: A Survey

### 12 Dec 2023 v1

@@ 1 Introduction
Specifically, our categorization of mathematical tasks (§2) contain two primary domains: mathematical calculation (§2.1), consisting of arithmetic representation and arithmetic calculation, and mathematical reasoning (§2.2), consisting of problem-solving and theorem proving. Furthermore, our taxonomy of existing algorithms segregates them into PLMs-based approaches (§3), including autoregression (§3.1) and non-autoregression (§3.2) LMs, and LLMs-based methodologies (§4), which include instruction learning (§4.1), tool-based strategies (§4.2), fundamental CoT techniques (§4.3), and advanced CoT methodologies (§4.4). Moreover, we provide a comprehensive compilation of over 60 mathematical datasets (§5), systematically categorized into training (§5.1), benchmark (§5.2), and augmented datasets (§5.3). This classification aims to facilitate research by delineating the utility and applicability of these datasets within distinct research contexts. Lastly, our survey meticulously addresses the primary challenges and explores the future directions of this domain (§6), including faithfulness, multi-modal, uncertainty, evaluation, creation and application. In conclusion (§7), we believe that this comprehensive survey will serve as a valuable asset for researchers and practitioners alike, seeking to push the boundaries of innovation in the field of mathematical language models.
@@ 2.2.1 Math Problem Solving
Recent research has highlighted the growing capabilities of LLMs in the field of mathematical word problem solving, emphasizing the trend toward more nuanced and sophisticated AI-driven mathematical analysis. MathPrompter [imani2023mathprompter] uses GPT-3 Davinci [brown2020language] to solve MWPs with excellent results, demonstrating the potential of LLMs to not only explain but also generate complex mathematical reasoning, reflecting a human-like understanding of complex problem sets.
@@ 2.2.2 Theorem Proving
Ongoing research is delving into the integration of language models and interactive proof assistants within the realm of mathematics. This involves generating proofs through tactic prediction [DBLP:journals/corr/abs-2009-03393, han2021proof, lample2022hypertree, jiang2022thor], automating formalization processes [DBLP:conf/nips/WuJLRSJS22, jiang2022draft], and developing unified tools [llmstep]. Due to the substantial computational demands of exploration, language models applied in this domain have traditionally been limited in size. However, recent advancements have showcased potential in employing larger models [first2023baldur, jiang2022draft]. The work of Azerbayev et al. (LLEMMA) [azerbayev8llemma] demonstrates few-shot proof autoformalization and tactic prediction, offering a vast dataset of formal mathematics and an openly accessible model for further investigation in these directions. These endeavors establish crucial theoretical and practical frameworks for automating and enhancing tactic prediction.
@@ 3 PLMs-based Methods
Documentation [hsieh2023tool] , leaf, text width=44.6em ] ] ] [ Fundamental CoT
@@
Methods [ ScoREval [he2023socreval], Socratic Models [zeng2022socratic], SOCRATIC QUESTIONING [qi2023art] , leaf, text width=44.6em ] ] ] ] ]
@@ 3.1 Autoregression LMs
To enhance the performance of PLMs in mathematical problems, many improved methods based on ALMs have been proposed. GPT-$f$ [polu2020generative] presents an automated prover and proof assistant to explore the performance of LMs to automated theorem proving. To resolve challenges related to expression fragmentation and operand-context separation, Kim et al. [kim_point_2020] introduce the Expression-Pointer Transformer (EPT), a pure neural model that employs both an ‘Expression’ token and operand-context pointers during the generation of solution equations. Generate &\&& Rank [shen_generate_2021] introduces a novel ranking task for MWP, a multi-task framework built upon a generative PLM, effectively addressing the challenge of minor mistakes in mathematical expressions. Thor [jiang2022thor] introduces a framework that integrates language models and automated TP. The latter is employed to selectively choose relevant premises from an extensive library to assist the language model in TP. HTPS [lample_hypertree_nodate] presents HyperTree Proof Search, a novel algorithm leveraging online training from past proof searches, enhancing model performance and surpassing the previous SOTA GPT-$f$. Galactica [taylor_galactica_2022] proposes a working memory token approach that can achieve strong performance over existing methods on mathematical MMLU [hendryckstest2021] and MATH [hendrycks2021measuring] benchmarks.
@@ 4.1 Instruction Learning
A semi-supervised approach, as presented in [nakamoto2023enhancing], leverages LLMs to create datasets for automating the scoring of mathematical self-explanations in the context of mathematics education. A challenging benchmark ProofNet [azerbayev2023proofnet] is introduced to build a system of automated theorem proving. Meanwhile, prompt retrieval and distilled back translation methods are introduced for statement autoformalization. WizardLM [xu_wizardlm_2023] presents a groundbreaking method called Evol-Instruct, designed to autonomously generate high-quality instructions by LLMs themselves. The conceptual flow is depicted in Figure 4. The instructional process commences with a straightforward “initial instruct" and progresses into diverse forms, deepening, increase reasoning, add constraints, concretizing, complicate Input (formula), in-Breadth evolving, etc. Furthermore, Wizardmath [luo2023wizardmath] proposes a reinforced Evol-instruct method (RLEIF) to build more complexity of instruct dataset. This method effectively enhances the mathematical reasoning capabilities of Llama-2 by supervised fine-tuning and PPO training [schulman2017proximal].
@@ 4.2 Tool-based Methods
Furthermore, CRITIC [gou2023critic] significantly enhances the outputs of LLMs by enabling them to verify and self-correct through interactions with external tools. Inspired by human cognition and critical thinking, CRITIC continuously refines text generated by LLMs through iterative interactions with tools like search engines and code interpreters. Tool Augmented Language Models (TALM) [parisi2022talm] seamlessly integrate language models with non-differentiable tools. They employ an iterative “self-play" method to bootstrap performance based on a limited number of tool demonstrations. In contrast, Hsieh et al. [hsieh2023tool] opted for tools based on documentation rather than relying on demonstrations. Current practices involve teaching LLMs through a few-shot demonstration of tool usage, a process prone to bias and overfitting. The alternative proposed by Hsieh et al., utilizing tool documentation that provides descriptions of individual tool usage, is argued to be a more effective approach. ToRA [gou2023tora] employs a set of Tool-integrated Reasoning Agents, which integrate computational libraries and symbolic solvers, among other mathematical tools, to effectively solve complex mathematical problems.
@@ 4.3 Fundamental CoT Methods
Utilizing the in-context learning of LLMs, CoT [wei2022chain] explicitly introduces the concept of the chain of thought (CoT) for the first time, and substantiates its efficacy in enhancing reasoning abilities. A CoT example is shown in Figure 8. This validation is also corroborated in Complexity-based CoT [fu2022complexity]. Following a similar line of thought, LVSS [lightman2023lets] presents a Process Reward Model (PRM) and compares it with the Outcome Reward Model (ORM). Empirical findings affirm that process supervision outperforms outcome supervision, leading to notable performance improvements on the MATH [hendrycks2021measuring] benchmarks. Consequently, several methods [chen2022program, yuan2023scaling, yu2023thought, wang2023making, sel2023algorithm, yue2023mammoth] leverage CoT methodologies to enhance the mathematical performance of models.
@@ 4.4 Advanced CoT Methods
LLMs often produce incorrect reasoning steps, which can lead to a series of cascading errors in their solutions. Implementing verification and refining the reasoning process based on feedback can significantly reduce these errors. This approach is akin to human reflection and involves critically evaluating each step of the reasoning process. Various methods [zhou2023solving, ling2023deductive, li2023making, zhao2023verify, he_rethinking_2022, shridhar2023screws] have been proposed to address these issues.
@@
The Socratic teaching method [nelson1980socratic], rooted in the philosophies of ancient Greek thinker Socrates, offers a distinctive perspective when considering the integration of large-scale language models (LLMs) in mathematics education [davies2021advancing]. This approach, emphasizing questioning and dialogue, aims to lead learners toward deep thinking [brickhouse2009socratic]. In field of mathematics education, employing this method can aid students in grasping concepts with clarity, fostering more rigorous logical application, and promoting reflection and self-discovery.
When applied to LLMs like GPT-3, utilizing the Socratic teaching method initiates a guided dialogue between the LLMs and the user [davis1997dialogue]. According to Chang et al. [chang2023prompting], employing techniques such as definition, dialectics, cross-examination, induction, and counterfactual reasoning enables LLMs to gain a deeper understanding of user intentions, thereby assisting users more effectively in addressing queries. These methods not only enhance the model’s comprehension but also, when combined with traditional inductive, deductive, and analogical reasoning techniques, establish a comprehensive analytical framework for tackling mathematical problems [chang2023prompting, alhossami2023language].
The modular framework of Socratic Models (SMs) [zeng2022socratic] significantly broadens the scope of interactive capabilities. SMs facilitate the amalgamation of multiple LLMs to tackle novel multi-modal tasks using structured Socratic dialogues. This approach proves highly effective in integrating diverse inputs seamlessly without necessitating additional training. Expanding on this Socratic teaching, the SocREval method [he2023socreval] focuses on evaluating the performance of LLMs in complex mathematical reasoning tasks. By automatically evaluating the quality of reasoning chains, this method offers a novel reference-free approach to reasoning assessment. This not only reduces time and costs but also enhances the evaluation’s precision.
Insights in this domain are further illuminated by studies like Zeng et al. [zeng2022socratic] and Shridhar et al. [shridhar2022automatic]. Qi et al. [qi2023art] and its counterpart that explores zero-shot multi-modal reasoning. These works delve into how recursive thinking and self-questioning can augment the problem-solving capabilities of LLMs. Moreover, contributions from Ang et al. [ang2023socratic] and Al et al. [al2023socratic] provide invaluable datasets and evaluation methodologies for Socratic questioning. These contributions underscore the practical applications and potential impact of Socratic methods in counseling and coaching, showcasing their diverse applicability.
@@ 5 Datasets
To train and evaluate the arithmetic and math ability of language models, various math word problems datasets [saxton2018analysing, hendrycks2021measuring, cobbe2021training, kushman-etal-2014-learning] are designed. In this paper, we organize the frequently used datasets for mathematical language models by dividing them into training, benchmark and data augmentation datasets (Table 1), and also sort out the common evaluation metrics.
@@ 6 Challenges and Further Directions
Mathematical LLMs reside in the phenomena of hallucination and faithfulness, where the generated output may lack factual accuracy or logical grounding [rawte2023survey, ji2023survey]. This phenomenon leads to the production of erroneous or misleading mathematical results, undermining the reliability of the model’s outputs. Some studies explored to address this problem by integrating extra knowledge [he_rethinking_2022], reinforcement learning from human feedback [lightman2023lets], tools [schick2023toolformer, parisi2022talm]. However, the improvement of the hallucination phenomena is limited, which influences the trustworthiness and utility of mathematical language models in practical applications and scholarly pursuits.
@@ 7 Conclusions
The survey elucidates the pivotal role of mathematical LMs in reshaping the landscape of mathematical problem-solving, leveraging a spectrum of models, from pre-trained language models (PLMs) to large-scale language models (LLMs), to address diverse mathematical tasks. Our taxonomical delineation of mathematical tasks and algorithms provides a systematic framework for comprehending the intricacies of LMs-based methodologies, distinguishing between arithmetic calculation, mathematical reasoning, and various algorithmic approaches employed in these models. The compilation of over 60 diverse mathematical datasets, categorized meticulously into training, benchmark, and augmented datasets, underscores the pivotal role of data in advancing mathematical research, facilitating informed research endeavors within distinct mathematical contexts. Moreover, by critically addressing challenges such as faithfulness, multi-modality, uncertainty, evaluation, theorem creation, and application, this survey paves the way for future investigations aimed at refining and advancing mathematical LMs capabilities. By shedding light on the current state-of-the-art, challenges, and avenues for future exploration, we envision this comprehensive overview to be a cornerstone in driving innovation and shaping the trajectory of mathematical LMs research, ultimately contributing to the evolving landscape of mathematics and artificial intelligence.

### 14 Dec 2023 v2

@@ 1 Introduction
Specifically, our categorization of mathematical tasks (§2) contain two primary domains: mathematical calculation (§2.1), consisting of arithmetic representation and arithmetic calculation, and mathematical reasoning (§2.2), consisting of problem-solving and theorem proving. Furthermore, our taxonomy of existing algorithms segregates them into PLMs-based approaches (§3), including autoregression (§3.1) and non-autoregression (§3.2) LMs, and LLMs-based methodologies (§4), which include instruction learning (§4.1), tool-based strategies (§4.2), fundamental CoT techniques (§4.3), and advanced CoT methodologies (§4.4). Moreover, we provide a comprehensive compilation of over 60 mathematical datasets (§5), systematically categorized into training (§5.1), benchmark (§5.2), and augmented datasets (§5.3). This classification aims to facilitate research by delineating the utility and applicability of these datasets within distinct research contexts. Lastly, our survey meticulously addresses the primary challenges and explores the future directions of this domain (§6), including faithfulness, multi-modal, uncertainty, evaluation, creation, application, and data scarcity. In conclusion (§7), we believe that this comprehensive survey will serve as a valuable asset for researchers and practitioners alike, seeking to push the boundaries of innovation in the field of mathematical language models.
@@ 2.2.1 Math Problem Solving
Recent research has highlighted the growing capabilities of LLMs in the field of mathematical word problem solving, emphasizing the trend toward more nuanced and sophisticated AI-driven mathematical analysis. MathPrompter [imani2023mathprompter] uses LLM called GPT3 DaVinci to solve MWPs with excellent results, demonstrating the potential of LLMs to not only explain but also generate complex mathematical reasoning, reflecting a human-like understanding of complex problem sets.
@@ 2.2.2 Theorem Proving
Ongoing research is delving into the integration of language models and interactive proof assistants within the realm of mathematics. This involves generating proofs through tactic prediction [DBLP:journals/corr/abs-2009-03393, han2021proof, lample2022hypertree, jiang2022thor], automating formalization processes [DBLP:conf/nips/WuJLRSJS22, jiang2022draft], and developing unified tools [llmstep]. Due to the substantial computational demands of exploration, language models applied in this domain have traditionally been limited in size. However, recent advancements have showcased potential in employing larger models [first2023baldur, jiang2022draft]. LLEMMA [azerbayev8llemma] demonstrates few-shot proof autoformalization and tactic prediction, offering a vast dataset of formal mathematics and an openly accessible model for further investigation in these directions. These endeavors establish crucial theoretical and practical frameworks for automating and enhancing tactic prediction.
@@ 3 PLMs-based Methods
Tool-Documentation [hsieh2023tool] , leaf, text width=44.6em ] ] ] [ Fundamental CoT
@@
Methods [ Socratic Models [zeng2022socratic], SOCRATIC QUESTIONING [qi2023art], Socratic prompt [chang2023prompting], SocratiQ [al2023socratic]
multi-turn Socratic advice [ang2023socratic] , leaf, text width=44.6em ] ] ] ] ]
@@ 3.1 Autoregression LMs
To enhance the performance of PLMs in mathematical problems, many improved methods based on ALMs have been proposed. GPT-$f$ [polu2020generative] presents an automated prover and proof assistant to explore the performance of LMs to automated theorem proving. To resolve challenges related to expression fragmentation and operand-context separation, Kim et al. [kim_point_2020] introduce the Expression-Pointer Transformer (EPT), a pure neural model that employs both an ‘Expression’ token and operand-context pointers during the generation of solution equations. Generate &\&& Rank [shen_generate_2021] introduces a novel ranking task for MWPs, a multi-task framework built upon a generative PLM, effectively addressing the challenge of minor mistakes in mathematical expressions. Thor [jiang2022thor] introduces a framework that integrates language models and automated TP. The latter is employed to selectively choose relevant premises from an extensive library to assist the language model in TP. HTPS [lample_hypertree_nodate] presents HyperTree Proof Search, a novel algorithm leveraging online training from past proof searches, enhancing model performance and surpassing the previous SOTA GPT-$f$. Galactica [taylor_galactica_2022] proposes a working memory token approach that can achieve strong performance over existing methods on mathematical MMLU [hendryckstest2021] and MATH [hendrycks2021measuring] benchmarks.
@@ 4.1 Instruction Learning
A semi-supervised approach, as presented in Auto-explanation [nakamoto2023enhancing], leverages LLMs to create datasets for automating the scoring of mathematical self-explanations in the context of mathematics education. A challenging benchmark ProofNet [azerbayev2023proofnet] is introduced to build a system of automated theorem proving. Meanwhile, prompt retrieval and distilled back translation methods are introduced for statement autoformalization. WizardLM [xu_wizardlm_2023] presents a groundbreaking method called Evol-Instruct, designed to autonomously generate high-quality instructions by LLMs themselves. The conceptual flow is depicted in Figure 4. The instructional process commences with a straightforward “initial instruct" and progresses into diverse forms, deepening, increase reasoning, add constraints, concretizing, complicate Input (formula), in-Breadth evolving, etc. Furthermore, Wizardmath [luo2023wizardmath] proposes a reinforced Evol-instruct method (RLEIF) to build more complexity of instruct dataset. This method effectively enhances the mathematical reasoning capabilities of Llama-2 by supervised fine-tuning and PPO training [schulman2017proximal].
@@ 4.2 Tool-based Methods
Furthermore, CRITIC [gou2023critic] significantly enhances the outputs of LLMs by enabling them to verify and self-correct through interactions with external tools. Inspired by human cognition and critical thinking, CRITIC continuously refines text generated by LLMs through iterative interactions with tools like search engines and code interpreters. Tool Augmented Language Models (TALM) [parisi2022talm] seamlessly integrate language models with non-differentiable tools. They employ an iterative “self-play" method to bootstrap performance based on a limited number of tool demonstrations. In contrast, Tool-Documentation [hsieh2023tool] opted for tools based on documentation rather than relying on demonstrations. Current practices involve teaching LLMs through a few-shot demonstration of tool usage, a process prone to bias and overfitting. The alternative proposed by Tool-Documentation, utilizing tool documentation that provides descriptions of individual tool usage, is argued to be a more effective approach. ToRA [gou2023tora] employs a set of Tool-integrated Reasoning Agents, which integrate computational libraries and symbolic solvers, among other mathematical tools, to effectively solve complex mathematical problems.
@@ 4.3 Fundamental CoT Methods
Utilizing the in-context learning of LLMs, CoT [wei2022chain] explicitly introduces the concept of the chain of thought (CoT) for the first time, and substantiates its efficacy in enhancing reasoning abilities. A CoT example is shown in Figure 8, where the text highlighted in blue and green guides to the ultimately correct answer. This validation is also corroborated in Complexity-based CoT [fu2022complexity]. Following a similar line of thought, LVSS [lightman2023lets] presents a Process Reward Model (PRM) and compares it with the Outcome Reward Model (ORM). Empirical findings affirm that process supervision outperforms outcome supervision, leading to notable performance improvements on the MATH [hendrycks2021measuring] benchmarks. Consequently, several methods [chen2022program, yuan2023scaling, yu2023thought, wang2023making, sel2023algorithm, yue2023mammoth] leverage CoT methodologies to enhance the mathematical performance of models.
@@ 4.4 Advanced CoT Methods
LLMs often produce incorrect reasoning steps, which can lead to a series of cascading errors in their solutions. Implementing verification and refining the reasoning process based on feedback can significantly reduce these errors. This approach is akin to human reflection and involves critically evaluating each step of the reasoning process. Various verify-based methods [zhou2023solving, ling2023deductive, li2023making, zhao2023verify, he_rethinking_2022, shridhar2023screws] have been proposed to address these issues.
@@
The Socratic teaching method [nelson1980socratic], derived from the philosophies of ancient Greek thinker Socrates, provides a unique perspective when considering the integration of LMs in mathematics education [davies2021advancing]. This method, emphasizing questioning and dialogue, aims to guide learners toward profound thinking [davis1997dialogue, brickhouse2009socratic]. In realm of mathematics education, employing this method can aid students in grasping concepts with clarity, fostering more rigorous logical application, and promoting reflection and self-discovery.
For LLMs, research shows that the Socratic teaching method can also improve their abilities to a certain extent. For example, the framework of Socratic Models (SMs) [zeng2022socratic] significantly broadens the scope of interactive capabilities. SMs facilitate the amalgamation of multiple LLMs to tackle novel multi-modal tasks using structured Socratic dialogues. In Chang et al. [chang2023prompting], techniques including definition, dialectics, cross-examination, induction, and counterfactual reasoning are utilized. These techniques contribute to enhancing the LLMs’ comprehension of user intentions, thereby improving model’s effectiveness in addressing user queries. The inductive, deductive, and analogical reasoning techniques in the Socratic teaching method establish a comprehensive analytical framework for solving mathematical problems [chang2023prompting, alhossami2023language], ensuring a strong and concise ability to address the complexities of mathematical problem-solving. Moreover, contributions from Ang et al. [ang2023socratic] and Al et al. [al2023socratic] provided invaluable datasets and evaluation methodologies for Socratic questioning. These contributions underscore the practical applications and potential impact of Socratic methods in counseling and coaching, showcasing their diverse applicability.
@@ 5 Datasets
To train and evaluate the arithmetic and math ability of language models, various math word problems datasets [saxton2018analysing, hendrycks2021measuring, cobbe2021training, kushman-etal-2014-learning] are designed. In this paper, we organize the frequently used datasets for mathematical language models by dividing them into training, benchmark and data augmentation datasets (Table 1).
@@ 6 Challenges and Further Directions
Mathematical LLMs reside in the phenomena of hallucination and faithfulness, where the generated output may lack factual accuracy or logical grounding [rawte2023survey, ji2023survey]. This phenomenon leads to the production of erroneous or misleading mathematical results, undermining the reliability of the model’s outputs. Some studies explored to address this problem by integrating extra knowledge [he_rethinking_2022], reinforcement learning from human feedback [lightman2023lets], tools [schick2023toolformer, parisi2022talm], and verify-based methods [zhou2023solving, ling2023deductive, li2023making, zhao2023verify, he_rethinking_2022, shridhar2023screws]. However, the improvement of the hallucination phenomena is limited, which influences the trustworthiness and utility of mathematical language models in practical applications and scholarly pursuits.
@@
Data scarcity.
The training data significantly influences the performance of language models, particularly in LLMs [ding_enhancing_2023]. High-quality and diverse training data can assist the model in enhancing its mathematical reasoning capabilities [luo2023wizardmath]. As discussed in Section 4.1, while there have been limited studies on constructing instruction data through LLMs, these efforts have only considered building from a small set of mathematical reasoning datasets, such as GSM8k [cobbe2021training] and MATH [hendrycks2021measuring]. High-quality and diverse mathematical instruction data remains scarce. There is a necessity to explore and investigate additional forms and construction methods for mathematical train data. Additionally, the generation of mathematical training datasets in a multimodal context is also a promising direction.
@@ 7 Conclusions
The survey elucidates the pivotal role of mathematical LMs in reshaping the landscape of mathematical problem-solving, leveraging a spectrum of models, from pre-trained language models (PLMs) to large-scale language models (LLMs), to address diverse mathematical tasks. Our taxonomical delineation of mathematical tasks and methods provides a systematic framework for comprehending the intricacies of LMs-based methodologies, distinguishing between arithmetic calculation, mathematical reasoning, and various algorithmic approaches employed in these models. The compilation of over 60 diverse mathematical datasets, categorized meticulously into training, benchmark, and augmented datasets, underscores the pivotal role of data in advancing mathematical research, facilitating informed research endeavors within distinct mathematical contexts. Moreover, by critically addressing challenges such as faithfulness, multi-modality, uncertainty, evaluation, theorem creation, application, and Data scarcity, this survey paves the way for future investigations aimed at refining and advancing mathematical LMs capabilities. By shedding light on the current state-of-the-art, challenges, and avenues for future exploration, we envision this comprehensive overview to be a cornerstone in driving innovation and shaping the trajectory of mathematical LMs research, ultimately contributing to the evolving landscape of mathematics and artificial intelligence.